4. The first product took 10 hours to build and the learning curve is estimated to be 90%. How long will it take to make the eighth unit?


TN = T1C


5. The first unit of production took 10 hours with a 90% learning curve. How long will it take to complete both the first and second units?


TN = T1C




1
Expert's answer
2021-01-24T18:43:14-0500

"T_1\times L^n=T_n," time required for the nth unit

where "T_1=" unit time of the first unit

"L=" learning curve raten

"n=" number of times T is doubled

4.

Hours required for unit 8 "=10\times(0.9)^3=7.29" hours.


5.

Hours required for unit 2 "=10\times(0.9)^1=9" hours.


"10" hours "+9" hours"=19" hours

It takes 19 hours to complete both the first and second units.